OpenStudy (anonymous):

A toy car with a mass of 2.5 kg travels at a velocity of 37 m/s for 2.5 s until it is stopped by a wall. What was the average acceleration of the toy car?

OpenStudy (anonymous):

@douglaswinslowcooper

OpenStudy (anonymous):

A. 14.8 m/s² B. 0.07 m/s² C. 14.8 m/s D. 0.07 m/s

OpenStudy (anonymous):

Units for acceleration are?

OpenStudy (anonymous):

Units?

OpenStudy (anonymous):

Acceleration = change in velocity / time = 37/2.5

OpenStudy (anonymous):

"units" means "dimensions" like kg, m, s etc.

OpenStudy (anonymous):

So 14.8?

OpenStudy (anonymous):

Yes, but I want you to now the units for acceleration velocity is length/time = m/s acceleration is velocity/time = (m/s)/s acceleration = ?

OpenStudy (anonymous):

So its C

OpenStudy (anonymous):

acceleration is m/s^2

hartnn (hartnn):

the formula is v = u+at u = initial vel = 37 v = final vel = 0 t = time = 2.5 a = acceleration

OpenStudy (anonymous):

Oh :( its A than

OpenStudy (anonymous):

Just so. Are we done? Have a good rest.

OpenStudy (anonymous):

I suck so much at these things :(

OpenStudy (anonymous):

Yes! Thanks @douglaswinslowcooper You helped a lot